Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report Page - 20121016

Observers : Patrick El-Hage, Assaf Sternberg. Data reduction team : Nancy and Antonia

All times are in UT

23:30 : SOFI calibration obs using Hip12852
00:00 : SOFI followup LSQ12dwl RG spec
02:10 : SOFI calibration obs using Hpi110512
02:20 : SOFI followup LSQ12dwl BG spec
04:45 : SOFI calibration obs using Hip110512
05:00 : SOFI followup SN2009ip BG, RG spec
05:45 : Conditions cloudy
06:15 : EFOSC followup LSQ12fhs gm11,gm16
06:35 : Lost guiding due to clouds, abandoning LSQ12fhs
06:35 : EFOSC followup CSS121008:014245 213928 gm13
07:20 : EFOSC followup SNhunt121 VRI phot
08:10 : LSQ12fiq classification
09:00 : Cloud cover preventing us from observing any calibration star as all EFOSC standards are too low at this hour.
Comments